OpenStudy (anonymous):

(-x^2y^4)(-x^3y^10)/x^9y

OpenStudy (angela210793):

=x^5y^14/x^9y=x^5y13

OpenStudy (saifoo.khan):

angela got that right

OpenStudy (anonymous):

Thanks, you know that the 2 equations in parenthesis are over the x^9 and y right?

OpenStudy (angela210793):

yes i know :)

OpenStudy (anonymous):

Thanks, you know that the 2 equations in parenthesis are over the x^9 and y right?

OpenStudy (anonymous):

lol okay thanks (: by the way you know how to make the problems look better like how to make it look like a fraction ? Any tools on here?? Thanks (:

OpenStudy (angela210793):

yea u cane use eq button =\[x^5y^13\]